Skip to content

OSG Technology Area Meeting, 2 July 2018

Coordinates: Conference: 719-284-5267, PIN: 57363; https://www.uberconference.com/osgblin
Attending: BrianL, Carl Derek, Marian, Mat, Suchandra, TimT

Announcements

  • CHEP (July 7 - July 13): Edgar attending
  • OSG User School (July 9 - July 13): BrianL, Derek, TimT attending

Triage Duty

  • This week: TimT
  • Next week: Mat
  • 12 (+5) open tickets

JIRA

# of tickets Δ State
125 +9 Open
17 -11 In Progress
10 -5 Ready for Testing
17 +15 Ready for Release

OSG Software Team

  • HDFS packages need to be untagged from upcoming after the release
  • New version of Singularity needs VMU tests kicked off
  • Planning the next doc focus: expect a doodle poll this week

Discussion

  • We will continue to ship Singularity even though Dave/Brian are EPEL maintainers due to yum priorities
  • WBS meeting yet to happen but once the work is done, it'll be shared with the Technology area teams
  • OASIS RPMs in goc-itb will be promoted to release tomorrow

Support Update

  • Topology (BrianL, Mat): Miscellaneous topology and downtime updates
  • University of Florida (Derek, Carl): We will modify the Slurm probe instead of trying out the HTCondor-CE probe

OSG Release Team

3.4.14 Δ Status
0 -6 Open
0 -7 In Progress
7 -6 Ready for Testing
17 +15 Ready for Release
24 -4 Total
  • OSG 3.4.14 (Release today)
    • Testing
      • RSV
        • fix cacert-verify-probe and crl-freshness-probe
      • An OSG lcmaps-plugins-voms patch breaks its use with llrun
      • Cleanup osg-se-hadoop-gridftp package
      • gratia
        • PBS/LSF update last timestamp for empty log file
        • htcondor-ce spam control
    • Ready for Release
      • GlideinWMS 3.4 (Upcoming)
      • OWAMP 3.5.6
      • RSV with Apache 2.4
      • gratia
        • twiki links
      • BLAHP - new version
      • HDFS 2.6
      • lcmaps-plugins-verify-proxy
      • osg-build
      • osg-configure
  • Data (Thursday release)
    • IGTF 1.92
  • GOC
    • Testing
      • Drop mirror.batlab.org from mirror list
    • Ready for Release
      • Nothing yet

Discussion

None this week

OSG Investigations Team

Operations

  • (Marian) Administration of OASIS
  • Moving OASIS to new hosts

TODO:

  • (Ongoing) Getting sites to transition to the new HTCondor collector
  • Discussion of SLA's, availability metrics
  • Puppetize: Collectors, GRACC

Investigations

  • Released StashCache client in production.

In Progress:

  • More StashCaches.
  • Beginning work with Internet2 to place StashCache in connection points.

Ongoing

Discussions

  • Anvil Cloud downtime: HCC offered hardware to hold us over and they should be up 1 week before the downtime. The exact downtime dates aren't scheduled yet.
  • Derek (AI): separate actual hosts for ITB host (instead of VHosts). Waiting on puppet modules.
  • Marian (AI): replica needs to be rebuilt, email BrianL, Derek, Jeff, TimC, and TimT about SLA and notification